# Privacy-protected Retrieval-Augmented Generation for Knowledge Graph Question Answering

### v1

@@ 4.3 Performance Comparison
Table 1: Comparison results (%) of Pure-LLM methods (Pure), SP-based methods (SP), and RAG-based methods (RAG), under the #Total setting (#Tot) and #Filtered setting (#Fil). The “-” indicates that the SP-based method cannot be run due to lacking special formal query annotations. “Imp.”: the absolute improvement of ARoG relative to the second-best performer. Bold: the best score. Underline: the second best score.
@@
We also conduct experiments on Qwen3-32B-FP8. The experiments yield findings that are consistent with the aforementioned conclusions. The details are shown in Appendix B.
@@ 4.4 Ablation Study
To assess the effectiveness of the two abstraction strategies in ARoG, we perform ablation studies on three datasets. Specifically, ablation studies are conducted separately on the three steps of the Relation-centric Abstraction module and on the Structure-oriented Abstraction module. Notably, relation retrieval (Step 1) aims to retrieve relations and serves as the foundation for all subsequent steps; relation filtering (Step 2) is a prerequisite for entity abstraction (Step 3). Therefore, in the ablation studies, we do not consider the scenarios of removing Step 1 or removing Step 3 while retaining Step 2. The results are shown in Table 2.
@@
Table 2: Results (%) of the ablation study. RA: Relation-centric Abstraction module. SA: Structure-oriented Abstraction module. #Tot: the #Total setting. #Fil: #Filtered setting. Bold: the best score.
@@ 4.5 Efficiency Study
We study the efficiency of RAG-based methods. Table 3 reports the average LLM calls and token consumption by these methods to answer a question across three datasets under the #Total setting.
Dataset Method #Call #Input #Output #Total
WebQSP ToG 15.4 6,356.4 1,357.0 7,713.4
GoG 10.6 12,031.0 343.3 12,374.3
PoG 10.1 5,872.5 336.0 6,208.6
ARoG 17.3 7,044.7 707.3 7,752.1
CWQ ToG 21.7 8,372.3 1,852.2 10,224.5
GoG 13.5 14,657.0 463.8 15,120.8
PoG 16.9 8,730.6 423.2 9,153.7
ARoG 25.4 1,0187.0 1,055.0 11,242.1
GrailQA ToG 13.5 5,220.1 1,306.5 6,526.5
GoG 11.4 12,665.7 428.0 13,093.7
PoG 15.8 7,445.0 436.4 7,881.4
ARoG 12.5 5,044.7 560.7 5,605.5
Table 3: Efficiency comparison between RAG-based methods under the #Total setting. #Call denotes the number of LLM calls, while #Input, #Output, #Total represent the number of input tokens, output tokens and total tokens, respectively.
On the WebQSP and CWQ datasets, we observe that ARoG is slightly less efficient than ToG and PoG. However, this disadvantage can be disregarded to some degree, especially considering ARoG’s outstanding performance. In contrast, the situation is different on the GrailQA dataset, whose most questions rely on external knowledge to answer. The reason is that ARoG can effectively utilize external knowledge while other RAG-based methods cannot, leading to higher consumption during the retrieval phase.
Additionally, we observe that GoG needs fewer LLM calls than most of others, but its input tokens consumption is too high. This is primarily due to its complex prompt design, which aims to decide whether to search knowledge from KGs or use LLMs to generate knowledge.
@@ 4.6 Quantitative Study
To understand the impact of retrieval depth $D$ and width $W$ on ARoG’s performance, we conduct the quantitative study. We vary these parameters from 1 to 4 (with a default of 3) across three datasets, following the approach in ToG (Sun et al., 2024). The results are presented in Figure 3.
@@ 4.7 Deep Analysis
To thoroughly validate the effectiveness of abstraction strategy in the Structure-oriented Abstraction module, we compare ARoG against two alternative abstraction strategies as follows. (1) Chain-of-Thought (CoT) (Wei et al., 2023), which replaces the abstract concept path with generated rationales, and (2) question decomposition (Dec) (Chen et al., 2024), which uses derived sub-questions. For a more comprehensive baseline, we also include the variant that omits the Structure-oriented Abstraction module (“w/o SA”). The performance of all these methods is illustrated in Figure 4.
@@
Case Study
The detailed case study is shown in the Appendix C.
@@ 5 Conclusion
Appendix
Appendix A Experiment Setup
A.1 Datasets
Table 4 details the dataset splits and the number of samples selected for our experiments. A notable distinction applies to the CWQ dataset, where the original annotations provide only single answer entity for each question. To facilitate a fair evaluation, we expand the answer set by executing the provided SPARQL (SPARQL Protocol and RDF Query Language) queries to retrieve all grounded entities. This methodology differs from that of ToG, GoG, and PoG, leading to a discrepancy in the reported performance on the CWQ dataset.
To ensure a fair comparison, both our proposed method and all baseline models are evaluated on the same knowledge graph. Consistent with the ToG framework, we utilize the Freebase public data dump (i.e. https://developers.google.com/freebase?hl=en), which we deploy using the Virtuoso RDF database system.
Dataset #(Train/Dev/Test) #Total #Filtered
WebQSP 3,098/-/1,639 1,639 535
CWQ 27,639/3,519/3,531 1,000 449
GrailQA 44,337/6,763/13,231 1,000 645
Table 4: Statistics of the datasets.
A.2 Settings for Baselines
IO, CoT and CoT-SC
For a fair comparison, we use the reproduced code for the IO, CoT, and CoT-SC methods from the ToG repository to infer answers.
KB-BINDER
We use the official repository and use KB-BINDER (6) (with majority vote) to infer answers.
TrustUQA
While we use the official repository for inference, we reconstruct the Condition Graph using the same knowledge graph as our method and other baselines to ensure a fair comparison. It is important to note that our reconstructed graph differs from the original in two significant ways: it is considerably larger in scale and exhibits a higher degree of noise.
ToG, GoG and PoG
For the baseline methods ToG, GoG, and PoG, we infer answers using their official repositories, adhering strictly to the default settings provided. To ensure a fair and consistent comparison, we evaluate all methods, including our proposed ARoG, using the evaluation protocol established in the official repository of ToG. Additionally, to investigate performance under a privacy-protected RAG scenario, we modify the Entity Name Search step, which prevents the leakage of sensitive entity information.
A.3 Implementation Details
In our method and all baselines, we employ gpt-4o-mini-2024-07-18 as the underlying LLM for a fair comparison, accessed via the OpenAI official API. For the Structure-oriented Abstraction module and the Generator model, the temperature parameter is set to 0 to ensure deterministic outputs. In contrast, for the Relation-centric Abstraction module, the temperature is set to 0.4 to introduce controlled variability. Both the frequency penalty and presence penalty are set to 0. The maximum token length for generation is capped at 256. Across all experiments, the width and depth of exploration are both set to 3, striking a balance between performance and efficiency. We repeat the experiment 3 times and report the average scores. All experiments are conducted on a server equipped with two Intel(R) Xeon(R) Silver 4310 CPU @ 2.10GHz and 128 GB RAM memory.
Appendix B Generalization Assessment
To validate the Generalization across different LLMs, we perform experiments using Qwen3-32B-FP8 to replace gpt-4o-mini-2024-07-18 for our method and all baselines. For a comprehensive evaluation, we benchmark ARoG against Pure-LLM and RAG-based baselines on the KGQA task. As detailed in Table 5, our method consistently outperforms these baselines.
The experiments yield findings that are consistent with the conclusions in the Section 4.3 Performance Comparison of main paper.
Type Method WebQSP CWQ GrailQA
#Tot #Fil #Tot #Fil #Tot #Fil
Pure IO 57.6 19.2 43.9 14.2 26.9 8.1
CoT 57.8 0.0 46.3 0.0 28.7 0.0
SC 57.7 3.3 46.1 1.3 29.1 2.4
SP KB-BINDER 20.2 9.1 - - 38.5 39.1
KB-BINDER-R 44.9 39.3 - - 60.5 60.9
TrustUQA 49.2 52.5 - - - -
TrustUQA-R 54.1 59.2 - - - -
RAG ToG 57.6 22.1 49.5 15.4 37.4 18.4
GoG 54.7 26.1 40.2 15.1 27.9 14.0
PoG 39.5 22.3 32.3 9.6 51.8 41.8
ARoG 72.7 64.3 54.3 36.5 78.7 74.8
Imp. $\uparrow$14.9 $\uparrow$5.1 $\uparrow$4.8 $\uparrow$21.1 $\uparrow$18.2 $\uparrow$13.9
Table 5: Comparison results (%) of Pure-LLM methods (Pure), SP-based methods (SP), and RAG-based methods (RAG), under the #Total setting (#Tot) and #Filtered setting (#Fil), using QWEN3-32B-FP8. The “-” indicates that the SP-based method cannot be run due to lacking special formal query annotations. “Imp.”: the absolute improvement of ARoG relative to the second-best performer. Bold: the best score. Underline: the second best score.
Appendix C Case Study
Figure 6 presents a typical case from the testing results on the CWQ dataset. We compare the results of ARoG, CoT, ToG, GoG and PoG in answering the question “What does the artist that was nominated for ‘The Audacity of Hope’ have a degree in?”. The underlying LLMs used are all GPT-4o-mini.
The case study illustrates the distinct operational mechanisms of each approach. In the first iteration of ARoG, it retrieves a set of question-relevant relations, such as “book.written_work.author”. It then identifies the corresponding entity cluster, which includes “m.02mjmr”. For this entity, ARoG searches and filters its adjacent relations, and then utilize an LLM to abstract it into “m.02mjmr (person)”. Concurrently, the Structure-oriented Abstraction module abstracts the question into the abstract concept path: “Common (artist) $\rightarrow$ nominated for $\rightarrow$ The Audacity of Hope (work); Common (artist) $\rightarrow$ has degree in $\rightarrow$ Communications (field of study)”. After that, the Abstraction-driven Retrieval module retrieves the question-relevant triplets, such as “The Audacity of Hope, book.written_work.author, m.02mjmr (person)”. All of the retrieved triplets are provided to the Generator module, and it generates a negative flag. Then, the topic entity is updated to “m.02mjmr (person)”. Starting from this new topic entity, the Abstraction-driven Retrieval module successfully retrieves “m.02mjmr (person), people.person.education-education.education.major_field_of_study, m.062z7 (academic discipline)”. Finally, the Generator module infers the correct answer “m.062z7 (academic discipline)”, i.e. Political Science.
In contrast, CoT relies on the LLM’s internal knowledge, which is incorrect in this case, it produces a wrong answer. Besides, ToG fails to identify the relevant anonymous entity “m.062z7” and instead retrieves incorrect triplets “m.02mjmr, award.award_nominee.award_nominations-award.award_nomination.award, m.0nh4p7s”. This failure occurs because ToG can not distinguish the entity “m.062z7” from other entities, such as “m.0nh4p7s”. Ultimately, ToG can not infer an answer and defaults to the result from the CoT approach. Similarly, although GoG and PoG introduce LLMs’ internal knowledge to supplement or modify the retrieved triplets, they introduce errors.
From this analysis, it is evident that ARoG outperforms CoT, ToG, GoG and PoG. ARoG success in retrieving question-relevant triplets and infer the correct answers, utilizing two abstraction strategies.
Figure 6: A typical case to compare different methods to answer the complex question under the privacy-protected RAG scenario for KGQA. Incorrect entities and wrong answers are highlighted in red. The noun concept of entities in the abstract concept path and retrieved triplets are highlighted in blue (question-relevant entities) or green (answer entities).
Appendix D Prompts
Here, we provide all the prompts used in ARoG. The specific in-context few-shot is shown in code files.
D.1 Relation-centric Abstraction
% The prompt used in the relation retrieval.
Please retrieve %s relations that contribute to the question from the following relations (separated by semicolons).
Q: Name the president of the country whose main spoken language was Brahui in 1980?
Topic Entity: Brahui Language
Relations: language.human_language.main_country; language.human_language.language_family; language.human_language.iso_639_3_code; base.rosetta.languoid.parent; language.human_language.writing_system; base.rosetta.languoid.languoid_class; language.human_language.countries_spoken_in; kg.object_profile.prominent_type; base.rosetta.languoid.document; base.ontologies.ontology_instance.equivalent_instances; base.rosetta.languoid.local_name; language.human_language.region
The output is:
[’language.human_language.main_country’,’language.human_language.countries_spoken_in’,’base.rosetta.languoid.parent’]
Q: {}
Topic Entity: {}
Relations: {}
% The prompt used in the entity abstraction. The ‘‘{{OBJECT_RELATIONS}}’’ and ‘‘{{SUBJECT_RELATIONS}}’’ are replaced by the related relations generated by the relation filtering.
Please provide the type with a short description of the ENTITY based on the context.
Here is an example:
Context:
ENTITY is the object of verbs: City/Town
ENTITY is the subject of verbs: Newspapers; Time zone(s); Area; Contained by; Population
The output is: {"type": "geographic location", "description": "a geographic location or administrative area that is associated with specific characteristics, such as a newspaper circulation areas, time zones, and population statistics."}
Context:
ENTITY is the object of verbs: {{OBJECT_RELATIONS}}
ENTITY is the subject of verbs: {{SUBJECT_RELATIONS}}
Now you need to directly output its type with a short description. Fill in the format below without other information or notes:
{"type": "the type of the entity", "description": "the short description of the entity"}
The output is:
D.2 Structure-oriented Abstraction
Please answer the questions step by step, the format of the reasoning process should be kept consistent with examples.
Question: What state is home to the university that is represented in sports by George Washington Colonials mens basketball?
Thought: First, the education institution has a sports team named George Washington Colonials mens basketball in is George Washington University. Second, George Washington University is in Washington D.C.. The answer is Washington, D.C..
Reasoning Path: {George Washington University (education institution) -> represented in sports by -> George Washington Colonials mens basketball (sport team)}; {George Washington University (education institution) -> located in -> Washington D.C. (state)}
Answer: Washington D.C. (state)
other in-context few-shot...
Question: {}
D.3 Generator
% The prompt used to output the flag.
Given a question and the associated retrieved knowledge graph triplets (subject entity, relation, object entity), you are asked to answer whether it is sufficient for you to answer the question with these triplets (Yes or No). Note that while several entities are represented as MIDs and corresponding type (e.g., Taylor Swift could be represented by m.0011 (person)’), you should assume the formal names are known.
Q: Find the person who said "Taste cannot be controlled by law", what did this person die from?
Knowledge Triplets: Taste cannot be controlled by law., media_common.quotation.author, m.0wfjf99 (person)
A: {No}. The retrieved knowledge graph triplet provides a quotation and identifies the author as a person (m.0wfjf99), but it does not provide any information about the cause of death of that person. To answer the question about what the person died from, additional information regarding the individuals death is required, which is not present in the provided triplet. Therefore, the triplet is insufficient to answer the question.
other in-context few-shot...
Q: {}
% The prompt used to output the answers.
Given a question and the associated retrieved knowledge graph triplets (entity, relation, entity), you are asked to answer the question with these triplets. Please note that while other entities are represented as IDs (e.g., m.0011’), you should assume you are familiar with their formal names.
Q: The artist nominated for The Long Winter lived where?
Knowledge Triplets: The Long Winter, book.written_work.author, m.0bvl_7 (person)
m.0bvl_7 (person), people.person.places_lived|people.place_lived.location, m.0wfjc51 (place)
A: {m.0wfjc51 (place)}. The artist nominated for The Long Winter, m.0bvl_7 (person), lived in a location that is represented by the entity m.0wfjc51.
other in-context few-shot...
Q: {}
Appendix E Search SPARQL
To automatically process the KG data in ARoG, we predefine the SPARQL for Freebase queries, which can be executed by filling in the entity’s mid and relation. Notably, we flatten Compound Value Type (CVT) nodes in Freebase into multiple single-hop relations in order to narrow the gap between candidate triplets and abstract concept paths.
E.1 Relations Search
% search all adjacent relations of a specific head entity.
PREFIX ns: <http://rdf.freebase.com/ns/>
SELECT ?relation
WHERE {
ns:%s ?relation ?x .
% search all adjacent relations of a specific tail entity.
PREFIX ns: <http://rdf.freebase.com/ns/>
SELECT ?relation
WHERE {
?x ?relation ns:%s .
E.2 Entity Cluster Search
% search all adjacent entity cluster of a specific head entity-relation pair.
PREFIX ns: <http://rdf.freebase.com/ns/>
SELECT ?tailEntity
WHERE {
ns:%s ns:%s ?tailEntity .
% search all adjacent entity cluster of a specific tail entity-relation pair.
PREFIX ns: <http://rdf.freebase.com/ns/>
SELECT ?tailEntity
WHERE {
?tailEntity ns:%s ns:%s .
% search all adjacent entity cluster of a specific head entity-relation pair, where the relation is flattened.
PREFIX ns: <http://rdf.freebase.com/ns/>
SELECT ?tailEntity
WHERE {
ns:%s ns:%s ?mid_entity .
?mid_entity ns:%s ?tailEntity .
% search all adjacent entity cluster of a specific tail entity-relation pair, where the relation is flattened.
PREFIX ns: <http://rdf.freebase.com/ns/>
SELECT ?tailEntity
WHERE {
?tailEntity ns:%s ?mid_entity .
?mid_entity ns:%s ns:%s .
E.3 Entity Name Search
PREFIX ns: <http://rdf.freebase.com/ns/>
SELECT DISTINCT ?tailEntity
WHERE {
FILTER (!isLiteral(?tailEntity) OR lang(?tailEntity) = ’’ OR langMatches(lang(?tailEntity), en’))
?entity ns:type.object.name ?tailEntity . FILTER(?entity = ns:%s)

### v2

@@ 4.3 Performance Comparison
Table 2: Comparison results (%) of Pure-LLM methods (Pure), SP-based methods (SP), and RAG-based methods (RAG), under the #Total setting (#Tot) and #Filtered setting (#Fil). “-”: the SP-based method cannot be run due to lacking special formal query annotations. “Imp.”: the absolute improvement of ARoG relative to the second-best performer. Bold: the best. Underline: the second best.
@@ 4.4 Ablation Study
To assess the effectiveness of the two abstraction strategies in ARoG, we perform ablation studies on three datasets. Specifically, ablation studies are conducted separately on the three steps of the Relation-centric Abstraction module and on the Structure-oriented Abstraction module. Notably, relation retrieval (Step 1) aims to retrieve relations and serves as the foundation for all subsequent steps; relation filtering (Step 2) acts as an enhancement to entity abstraction (Step 3). Therefore, in the ablation studies, we do not consider the scenarios of removing Step 1 or removing Step 3 while retaining Step 2. The results are shown in Table 3.
@@
Table 3: Results (%) of the ablation study. RA: Relation-centric Abstraction module. SA: Structure-oriented Abstraction module. #Tot: the #Total setting. #Fil: #Filtered setting.
@@ 4.5 Efficiency Study
We study the efficiency of RAG-based methods. Table 4 reports the average token usage and LLM calls under the #Total setting. In terms of total token usage, which directly reflects efficiency, ARoG achieves competitive or superior results. On WebQSP, its total token usage is close to ToG and much lower than GoG. On CWQ, ARoG maintains an affordable cost with fewer total token usage than GoG while keeping high performance. Notably, on the most challenging GrailQA dataset, ARoG achieves the lowest total token usage, demonstrating clear efficiency advantages in handling complex problems. Although ARoG requires more LLM calls in some cases, this metric is not directly related to efficiency, as the cost of individual calls varies.
Dataset Method #Input #Output #Total #Call
WebQSP ToG 6,356.4 1,357.0 7,713.4 15.4
GoG 12,031.0 343.3 12,374.3 10.6
PoG 5,872.5 336.0 6,208.6 10.1
ARoG 7,044.7 707.3 7,752.1 17.3
CWQ ToG 8,372.3 1,852.2 10,224.5 21.7
GoG 14,657.0 463.8 15,120.8 13.5
PoG 8,730.6 423.2 9,153.7 16.9
ARoG 10,187.0 1,055.0 11,242.1 25.4
GrailQA ToG 5,220.1 1,306.5 6,526.5 13.5
GoG 12,665.7 428.0 13,093.7 11.4
PoG 7,445.0 436.4 7,881.4 15.8
ARoG 5,044.7 560.7 5,605.5 12.5
Table 4: Efficiency comparison between RAG-based methods under the #Total setting. #Input, #Output, #Total represent the number of input tokens, output tokens, and total tokens, respectively. #Call denotes the number of LLM calls.
Overall, ARoG demonstrates strong token efficiency across datasets, indicating effective and privacy-protected reasoning with affordable computational cost.
@@ 4.6 Quantitative Study
To understand the impact of retrieval depth $D$ and width $W$ on ARoG’s performance, we conduct the quantitative study. We vary these parameters from 1 to 4 (with a default of 3) across three datasets, following the approach in ToG (sun2024thinkongraph). The results are presented in Figure 3.
@@ 4.7 Deep Analysis
To thoroughly validate the effectiveness of abstraction strategy in the Structure-oriented Abstraction module, we compare ARoG against two alternative abstraction strategies as follows. (1) Chain-of-Thought (CoT) (wei2023chainofthoughtpromptingelicitsreasoning), which replaces the abstract concept path with generated rationales, and (2) question decomposition (Dec) (NEURIPS_2024_POG), which uses derived sub-questions. For a more comprehensive baseline, we also include the variant that omits the Structure-oriented Abstraction module (“w/o SA”). The performance of all these methods is illustrated in Figure 4.
@@ 5 Conclusion
Acknowledgments
This work was supported by the National Natural Science Foundation of China (NSFC) (Grant No. 62576256), the Fundamental Research Funds for the Central Universities, China (Grant No. 2042022dx0001), and the Key Laboratory of Computing Power Network and Information Security, Ministry of Education (Grant No. 2024ZD027).